A flat circular disk of radius "a" is placed inside a uniform external electric field E' pointing in positive z direction. calculate the net electric flux through the disk.

1
Expert's answer
2022-01-09T18:22:30-0500

Assuming the plane of the disk is horizontal


"\phi=\int\>E.dA"


Since electric field is uniform


"\phi=EA\>Cos\>\theta"


"\theta=0°"


"\phi=EA"

But "A=\pi\>a^2"


"\phi=\pi\>Ea^2"


Where "E=" Electric field

"\phi=" Electric flux